# RLHFuse: Efficient RLHF Training for Large Language Models with Inter- and Intra-Stage Fusion

### v2

@@ 4.1 Opportunities and Challenges
Based on this, we design a fused execution plan that overlaps the generation and inference stages. Specifically, the fused plan first detects the point when the generation instance enters the long-tail decoding phase. At this juncture, most of the samples have finished their generation stage but are waiting for the inference stage to start. Each instance only leaves a few long-tailed samples to process, which is memory-bound as discussed in §2.2. This condition facilitates the migration of samples from all instances to a few designated instances. This migration not only consolidates the GPU utilization of the receiving instances but also releases resources from the sending instances. The released resources are then promptly utilized to launch inference tasks in advance. One example of a fused execution plan is shown in Figure 5. This approach greatly improves the GPU utilization and optimizes the overall execution time of the two stages.
@@ 4.2 Fused Execution Plan
To determine the optimal $R_{t}$, we initially conducted offline generation trials to analyze the length distribution of the generated samples. This analysis enables us to estimate the computational load for both the generation and inference stages accurately. We then systematically test $R_{t}$ values ranging from 5% to 95% of the global batch size, simulating the execution time for each value. The optimal $R_{t}$ is identified as the value that yields the minimum simulated execution time. Additionally, during runtime, we refine the distribution by incorporating new generation samples, which allows us to update $R_{t}$ as needed to continually optimize performance.
@@
Migration destination. Once the migration timing is determined, we need to select $m$ generation instances to handle the remaining samples. Assume there are $n$ generation instances in total, then there will be $n-m$ instances to be repurposed for inference tasks. To determine $m$, we revisit the computation characteristics of the generation task. During the memory-bound decoding phase of generation, as long as the batch size does not exceed a certain threshold that saturates the GPU (denoted as $BS_{max}$), the latency of each decoding iteration remains almost constant [26, 27]. The value of $BS_{max}$ depends on the specific GPU hardware and can be determined through prior profiling. Based on this, we set the first constraint as $m\geq\frac{R_{t}}{BS_{max}}$, which ensures that the generation time of long-tailed samples remain unchanged as before the migration.
@@
(a) Chimera                                                         (b) RLHFuse
Figure 6: (a) An example of the symmetric bi-directional pipeline schedule used in Chimera [20] with four pipeline stages and four micro-batchs. (b) An example to show our method can fuse two different models with four and two pipeline stages, respectively.
@@ 5.2 Fused Pipeline Schedule
Problem transformation. Assume the parallel strategies for models $A$ and $B$ are $(dp_{1},pp_{1},tp_{1})$ and $(dp_{2},pp_{2},tp_{2})$. Each will utilize all the GPUs in the cluster. We require that $tp$ is the powers of two, which is commonly adopted in practice. For the case where $tp_{1}\neq tp_{2}$, the challenge for a fused pipeline schedule lies in the fact that the pipeline stage of the two models contains a different number of GPUs. Without loss of generality, we assume $tp_{1}=s\times tp_{2}$. In this situation, we merge every $s$ consecutive pipeline stages of model $B$ into one stage, redividing model $B$ into $\frac{pp_{2}}{s}$ pipeline stages. This ensures that each stage of both models uses the same number of GPU resources. Note that here we require $pp_{2}$ to be divisible by $s$, which is easy to realize in practice. After that, we can transform the problem into fusing $K_{1}$ disjoint pipeline groups of model $A$ with $K_{2}$ pipeline groups of model $B$, where $K$ denotes the fusion factor and $K_{1}$ and $K_{2}$ are coprime. Figure 6(b) illustrates an example of $(K_{1},K_{2})=(1,2)$. Next, we precisely define the fused pipeline schedule and formulate the problem. Table 1 summarizes the key notations.
@@
Algorithm overview: The problem has many constraints and variables. It can be reduced to a special case of flow job scheduling problem which proves to be NP-hard [29]. A naive solution is to greedily fuse the 1F1B schedules of two models with the bi-directional pipeline, which always schedules feasible micro-batches. If the micro-batches of two models are both ready, it favors the larger model, with the expectation that the smaller one can flexibly fill in the bubbles later. However, the greedy method lacks a global perspective. For example, certain micro-batches could be deliberately delayed to fill the bubbles later. Moreover, it provides no optimization for memory usage. Thus, the greedy approach does not yield optimal performance, as we will demonstrate in §7.3.
@@
Simulated Annealing (Algorithm 1): The algorithm uses the solution found by the greedy algorithm as the initial state $S_{0}$, and its execution time as the initial temperature (line 2-3). $s^{*}$ is used to store the schedule with the lowest latency and $e^{*}$ is the energy (latency) of $s^{*}$. The algorithm searches until temperature $T$ is less than an epsilon value (line 7-16). $T$ is decreased by a factor of $\alpha$ in every iteration. At each iteration, it uses $ComputeNeighbor$ subroutine to find a neighbor state of the current one and uses $ComputeEnergy$ to compute the energy of the neighbor state. If the neighbor state has a lower energy than $s^{*}$, it updates $s^{*}$ (line 10-12). The algorithm uses a probabilistic function $P$ to decide whether to transition from the current state to the neighbor state. The probabilistic function $P$ is defined as follows: if the neighbor state has a lower energy than the current state, the probability is $1$; otherwise, the probability is $e^{(e_{current}-e_{neighbor})/T}$.
@@ 6 System Implementation and Optimizations
Parallelism strategy configuration. Optimizing the parallel strategy for LLM training is a well-studied problem [33, 16, 34, 13]. Due to the deterministic nature of LLM computation, the execution time and memory cost can be accurately modeled through minimal profiling. An optimal solution can then be found through optimizations [33, 25, 34, 13]. The generation and inference task involves only the forward pass, which is a subproblem of the LLM training. ReaLHF [13] is a state-of-the-art solution that optimizes the parallel strategy for RLHF training and RLHFuse adopts a similar model-then-optimize approach. We build a simulator to accurately estimate the runtime statistics of the tasks under a specific parallel strategy and workload. Then we follow the guidelines in the Megatron-LM [16] paper to prune the design space and brute-force search the optimal strategy.
@@
Stage fusion. For inter-stage fusion, we use the simulator mentioned above to choose the best migration threshold $R_{t}$ to minimize the fused execution time under a specific workload pattern. For intra-stage fusion, we use MPI [35] to parallelize the simulated annealing with different random seeds. After finding the optimal solution, it will generate the sequences of NCCL [36] operations on each device following the schedule. The runtime will follow the instruction flow to execute the fused pipeline schedule.
@@
System optimizations. Additionally, we adopt a series of system optimizations tailored for RLHF training. These optimizations are implemented in our own system as well as open-source RLHF frameworks like ReaLHF [13]. To separate the differences in low-level system implementations from the core techniques (inter-stage fusion in §4 and intra-stage fusion in §5), our evaluation in §7 uses RLHFuse-Base that includes these optimizations but without inter- and intra-stage fusion as an additional baseline.
Generation stage. We use an in-house inference engine with optimized CUDA kernels. It supports continuous batching [37] to allow the short samples to leave the batch early. For samples that share the same prompt, it supports prefix sharing [38] to avoid key-value cache duplication.
Inference stage. During the post-processing of the inference stage, one metric called Generalized Advantage Estimation [39] is needed which requires recursive computation along the output length dimension for each sample. When the output length is long, it incurs numerous kernel launches which takes even longer than the inference computation itself. To address this, we unroll the recursive formula along the output length dimension and precompute the coefficient matrix. This trick transforms the original recursive computation into a single matrix multiplication, successfully reducing the latency from tens of seconds to just tens of milliseconds.
Training stage. In RLHF training, DP groups need to synchronize after processing each mini-batch. Due to data skewness, different DP groups may have varying workloads, leading to stragglers and wasted GPU time. To mitigate this, we evenly distribute each mini-batch across the DP groups based on the sequence length of the samples, ensuring that the workloads are roughly balanced across the DP groups. Note that this will not violate the original training semantics.
Task switching. The primary overhead during task switching lies in model weights reloading. For the Actor and Critic models, the latest weights after the training stage need to be redistributed to the appropriate devices according to the new parallel strategy for generation or inference. This communication pattern is well-studied [40, 14, 13] and we adopt a similar all-to-all approach to minimize cross-node communication. For the Ref and RW models whose weights remain unchanged, we keep them in CPU memory and swap them into GPU memory as needed. The swapping can overlap with the computation of previous tasks. With these optimizations, the task switching overhead is negligible compared to the total iteration time, as shown in §7.2.
@@ 7 Evaluation
In this section, we evaluate RLHFuse under different sizes of LLMs ranging from 13B to 65B on real-world datasets. First, we compare the end-to-end performance of RLHFuse to other RLHF training frameworks (§7.1). Then we present the latency breakdown to delve into the performance improvement of RLHFuse (§7.2) and conduct ablation studies of our proposed techniques (§7.3). Finally, we provide a case study to show one execution plan generated by RLHFuse (§7.4).
Cluster testbed. We deploy RLHFuse on a production cluster for RLHF training with 32 nodes and 256 GPUs. Each node has 2TB of memory and 8 NVIDIA SXM H100-80GB GPUs connected with 900GB/s NVLINK. Nodes are connected by $8*200$ Gbps RDMA network based on RoCEv2 with rail-optimized topology. The simulated annealing algorithm in §5 is conducted on four CPU nodes, each with 2 AMD 9654 CPUs and there are 768 physical cores in total.
@@ 7.1 End-to-End Results
We compare the end-to-end performance of RLHFuse against the following RLHF training frameworks. For baselines, we explore all feasible configurable options and report the best performance achieved without out-of-memory errors.
  • DeepSpeed-Chat [44] (DSChat) only supports ZeRO-3 data parallelism [45] during training and utilizes a HybridEngine to switch from ZeRO-3 DP to TP in the generation stage. Since each GPU requires at least one sample during training and it only supports ZeRO-3 DP, we increase its mini-batch size to 256 while maintaining the original global batch size in order to run it successfully on our cluster. Note that with a larger batch size, this adjustment is more favorable for its throughput performance.
@@
  • RLHFuse-Base is RLHFuse without inter- and intra-stage fusion but with all the system optimizations enabled. We include this baseline to demonstrate the performance improvements brought by our core techniques in §4 and §5, eliminating any unfair comparisons caused by the differences in underlying framework implementations.
@@
Figure 9: The fused generation and inference stage execution time of RLHFuse under different migration ratio and model settings. The maximum generation length is set to 1024.
Figure 7 shows the end-to-end performance of the RLHF training systems under three generation settings. We set the maximum generation length to 512, 1024, and 2048, respectively. Under a specific generation setting, the configuration X/Y denotes parameter counts in actor and critic models, mirrored in the reference and reward models, respectively. Compared to DSChat, RLHFuse achieves 2.5×2.5$\times$2.5 ×3.7×3.7$\times$3.7 × higher throughput due to our flexible execution plan generator, which selects the most efficient 3D-parallel strategy for each task. Compared to ReaLHF, which also supports tailored parallel strategies, RLHFuse achieves 1.4×1.4$\times$1.4 ×2.4×2.4$\times$2.4 × higher throughput due to more fine-grained system optimizations. Specifically, we observe that the primary speedup comes from the customized kernels in the inference engine and the mini-batch load balancing during the training stage. Compared to RLHFuse-Base, our system achieves a relative improvement of 1.2×1.2$\times$1.2 ×1.4×1.4$\times$1.4 × on the throughput. This improvement is entirely from stage fusion, which significantly alleviates the issues of data skewness in the generation stage and pipeline bubbles in the training stage. In the following, we conduct a detailed breakdown analysis to show the performance improvement.
@@ 7.2 Performance Analysis
For Gen.+Inf., since RLHFuse-Base does not perform inter-stage fusion and executes the two stages in serial, we use light and dark colors to represent the time spent on the generation and inference stages respectively to better show the speedup of inter-stage fusing. It can be seen that as the maximum generation length increases, the processing time of long-tailed samples is long enough for RLHFuse to fully overlap the inference stage execution, achieving 1.2×1.2$\times$1.2 ×1.6×1.6$\times$1.6 × speedup. For the training stage, similarly, we use light and dark colors to represent the Actor and Critic training time in RLHFuse-Base. Through intra-stage fusion, RLHFuse greatly mitigates pipeline bubbles and reduces the execution time of the training stage by 1.2×1.2$\times$1.2 ×1.3×1.3$\times$1.3 ×. As for the other overhead, thanks to our optimizations in task switching (§6), it only accounts for less than 3% of the total iteration time, and the migration overhead of inter-stage fusion is also negligible.
Figure 10: The fused pipeline schedule of 65B Actor and 33B Critic model generated by RLHFuse. The two models have 16 and 8 PP stages respectively and #micro-batches is set equal to PP. Top: the GPU execution timeline. Bottom: the peak activation memory usage timeline. The dotted horizontal line denotes the memory cost of serial 1F1B execution.

### v3

@@ 4.1 Opportunities and Challenges
Based on this, we design a fused execution plan that overlaps the generation and inference stages. Specifically, the fused plan first detects the point when the generation instance enters the long-tail decoding phase. At this juncture, most of the samples have finished their generation stage but are waiting for the inference stage to start. Each instance only leaves a few long-tailed samples to process, which is memory-bandwidth-bound as discussed in §2.2. This condition facilitates the migration of samples from all instances to a few designated instances. This migration not only consolidates the GPU utilization of the receiving instances but also releases resources from the sending instances. The released resources are then promptly utilized to launch inference tasks in advance. One example of a fused execution plan is shown in Figure 5. This approach greatly improves the GPU utilization and optimizes the overall execution time of the two stages.
@@ 4.2 Fused Execution Plan
To determine the optimal $R_{t}$, we initially conducted offline generation trials on the training dataset to analyze the length distribution of the generated samples. This analysis enables us to estimate the computational load for both the generation and inference stages accurately and simulate the execution time of the fused execution plan. Due to the determinism and reproducibility of LLM computation, the simulation technique for LLM training [25, 26] and inference [27, 28] is well-studied. We then systematically test $R_{t}$ values ranging from 5% to 95% of the global batch size, simulating the overall execution time under each $R_{t}$. The optimal $R_{t}$ is identified as the value that yields the minimum simulated execution time. Additionally, during runtime, we refine the distribution by incorporating new generation samples, which allows us to update $R_{t}$ as needed to continually optimize performance.
@@
Migration destination. Once the migration timing is determined, we need to select $m$ generation instances to handle the remaining samples. Assume there are $n$ generation instances in total, then there will be $n-m$ instances to be repurposed for inference tasks. To determine $m$, we revisit the computation characteristics of the generation task. During the decoding phase of generation, which is highly memory-bandwidth-bound, as long as the batch size does not exceed a certain threshold that saturates the GPU (denoted as $BS_{max}$), the latency of each decoding iteration remains almost constant [28, 29] thanks to the massive parallel computing units of modern GPUs. The value of $BS_{max}$ depends on the specific GPU hardware and can be determined through prior profiling. Based on this, we set the first constraint as $m\geq\frac{R_{t}}{BS_{max}}$, which ensures that the generation time of long-tailed samples remain unchanged as before the migration.
@@
(a) Chimera [19]                                                  (b) RLHFuse
Figure 6: (a) An example of the symmetric bi-directional pipeline schedule used in Chimera [19] with four pipeline stages and four micro-batchs. (b) An example to show our method can fuse two different models with different numbers of pipeline stages.
@@ 5.2 Fused Pipeline Schedule
Problem transformation. Assume the parallel strategies for models $A$ and $B$ are $(dp_{1},pp_{1},tp_{1})$ and $(dp_{2},pp_{2},tp_{2})$, which denotes the parallel degree of each 3D-parallelism dimension. Each will utilize all the GPUs in the cluster. We require that $tp$ is the powers of two, which is commonly adopted in practice. For the case where $tp_{1}\neq tp_{2}$, the challenge for a fused pipeline schedule lies in the fact that the pipeline stage of the two models contains a different number of GPUs. Without loss of generality, we assume $tp_{1}=s\times tp_{2}$. In this situation, we merge every $s$ consecutive pipeline stages of model $B$ into one stage, redividing model $B$ into $\frac{pp_{2}}{s}$ pipeline stages. This ensures that each stage of both models uses the same number of GPU resources. Note that here we require $pp_{2}$ to be divisible by $s$, which is easy to realize in practice. After that, we can transform the problem into fusing $K_{1}$ pipeline groups of model $A$ with $K_{2}$ pipeline groups of model $B$, where $K_{i}$ denotes the fusion factor and $K_{1}$ and $K_{2}$ are coprime. Figure 6(b) illustrates an example of $(K_{1},K_{2})=(1,2)$. Next, we precisely define the fused pipeline schedule and formulate the problem. Table 1 summarizes the key notations.
@@
Algorithm overview: The problem has many constraints and variables. A naive solution is to extend the bi-directional pipeline [19] greedily which always schedules feasible micro-batches. If the micro-batches of two models are both ready, it favors the larger model, with the expectation that the smaller one can flexibly fill in the bubbles later. However, the greedy method lacks a global perspective. For example, certain micro-batches could be deliberately delayed to fill the bubbles later. Moreover, it provides no optimization for memory usage. Thus, the greedy approach does not yield optimal performance, as we will demonstrate in §7.3.
@@
Simulated Annealing (Algorithm 1): The algorithm uses the solution found by the greedy algorithm as the initial state $S_{0}$, and its execution time as the initial temperature (line 2-3). $s^{*}$ is used to store the schedule with the lowest latency and $e^{*}$ is the energy (latency) of $s^{*}$. The algorithm searches until temperature $T$ is less than an epsilon value (lines 7-16). $T$ is decreased by a factor of $\alpha$ in every iteration. At each iteration, it uses $ComputeNeighbor$ subroutine to find a neighbor state of the current one and uses $ComputeEnergy$ to compute the energy of the neighbor state. If the neighbor state has lower energy than $s^{*}$, it updates $s^{*}$ (lines 10-12). The algorithm uses a probabilistic function $P$ to decide whether to transition from the current state to the neighbor state. The probabilistic function $P$ is defined as follows: if the neighbor state has a lower energy than the current state, the probability is $1$; otherwise, the probability is $e^{(e_{current}-e_{neighbor})/T}$.
@@ 6 Implementation
Parallel strategy configuration. Optimizing the parallel strategy for LLM training task is a well-studied problem [26, 15, 34, 6]. Due to the deterministic nature of LLM computation [35], the execution time and memory cost can be accurately modeled through minimal profiling [25, 27, 28]. An optimal solution can then be found through optimizations [26, 18, 34, 6]. The generation and inference task involves only the forward pass, which can be modeled as a simplified problem of LLM training [27, 28]. ReaLHF [6] is a state-of-the-art solution that optimizes the parallel strategy for RLHF tasks and we adopt a similar model-then-optimize approach. We build a simulator to accurately estimate the runtime statistics of the RLHF task under a specific parallel strategy and workload pattern. Then we follow the guidelines in the Megatron-LM [15] paper to prune the design space and brute-force search the optimal strategy for each RLHF task.
Inter-stage fusion. We extend the simulator mentioned above from the task level to the workflow level. Given the parallel strategies, workload pattern, and migration threshold $R_{t}$, it considers the task dependencies and data transmission overhead and simulates the overall execution time of the generation and inference stages. Then we can choose the best migration threshold $R_{t}$ to minimize the fused execution time.
@@
Intra-stage fusion. We use MPI [36] to parallelize the simulated annealing algorithm used for fused pipeline schedule generation (§5.2). Under different random seeds, the computations are completely independent, so that we can easily scale the computation to hundreds of CPU cores, making it highly likely to find the optimal solution (§7.3). After finding the best solution, it will generate the sequences of NCCL [37] operations on each device following the fused pipeline schedule. The underlying execution engine will follow the instruction flow to execute the exact schedule during runtime. The code of the intra-stage fusion algorithm is publicly available [^1] https://github.com/FlexFusion/FlexFusion.
System optimizations. We adopt a series of system optimizations tailored for RLHF training. Since the RLHF training workflow is well-known, many of the following optimizations are also implemented or partially supported in a similar way in other RLHF frameworks like ReaLHF [6], HybridFlow [5], OpenRLHF [38], DeepSpeed-Chat [39], and PUZZLE [7].
For the generation stage, we implement an in-house inference engine which integrates most of the modern techniques optimized for LLM autoregressive generation like continuous batching [9], prefix sharing [10], and chunked-prefill [40]. For the inference stage, we optimize the computation for Generalized Advantage Estimation [41], which unrolls its original recursive formula along the output length dimension and transforms the recursive computation into a single matrix multiplication, greatly reducing the kernel launch overhead. For the training stage, we evenly distribute each mini-batch across the DP groups [6] based on the sequence length of the samples to ensure that the workloads are roughly balanced across the DP groups, effectively tackling the straggler issue. To reduce the task switching overhead [7, 5, 6], we minimize the cross-node communication [42] for the Actor and Critic models whose latest weights after the training stage need to be redistributed to the appropriate devices according to the new parallel strategies. For the Ref and RW models whose weights remain unchanged, we keep them in CPU memory and swap them into GPU memory as needed [5, 6], overlapping with the computation of previous tasks.
@@
To separate the differences in low-level system implementations from the core techniques (inter-stage fusion in §4 and intra-stage fusion in §5), our evaluation in §7 uses RLHFuse-Base that includes these optimizations but without inter- and intra-stage fusion as an additional baseline.
@@ 7 Evaluation
In this section, we evaluate RLHFuse under different sizes of LLMs ranging from 13B to 65B on real-world datasets.
Cluster testbed. We deploy RLHFuse on a production cluster for RLHF training with 32 nodes and 256 GPUs. Each node has 2TB of memory and 8 NVIDIA Hopper GPUs connected with NVLINK. Nodes are connected by $8*200$ Gbps RDMA network based on RoCEv2 with rail-optimized topology. The simulated annealing algorithm in §5 is conducted on four CPU nodes, each with 2 AMD 9654 CPUs and there are 768 physical cores in total.
@@ 7.1 End-to-End Results
We compare the end-to-end performance of RLHFuse against the following RLHF training frameworks.
  • DeepSpeed-Chat [39] (DSChat) colocates all the models on the same set of devices and only supports ZeRO-3 data parallelism [46] during training. It utilizes a HybridEngine to switch from ZeRO-3 DP to TP in the generation stage. Since each GPU requires at least one sample during training and it only supports ZeRO-3 DP, we increase its mini-batch size to 256 while maintaining the original global batch size in order to run it successfully on our testbed. Note that with a larger mini-batch size, this adjustment is more favorable for its throughput performance.
@@
  • RLHFuse-Base is RLHFuse without inter- and intra-stage fusion but with all the system optimizations enabled as described in §6. We include this baseline to demonstrate the performance improvements brought by our fusion techniques in §4 and §5, eliminating any unfair comparisons caused by the differences in underlying framework implementations and other optimization techniques in LLM generation and training which are orthogonal to the proposed stage fusion techniques.
@@
Figure 7 shows the end-to-end performance of the RLHF training systems under three generation settings. We set the maximum generation length to 512, 1024, and 2048, respectively. Under a specific generation setting, the configuration X/Y denotes parameter counts in actor and critic models, mirrored in the reference and reward models, respectively. Compared to DSChat, RLHFuse achieves 2.5×2.5$\times$2.5 ×3.7×3.7$\times$3.7 × higher throughput. This is because DeepSpeed-Chat colocates all the models on the same set of GPUs, making it impossible to adopt the most suitable parallel strategy for each task, which introduces additional computation and communication overhead. ReaLHF implements a flexible execution plan generator, which selects the most efficient 3D-parallel strategy for each task, such as balancing between DP and PP sizes to reduce the number of pipeline stages while avoiding OOM, thereby partially mitigating pipeline bubbles. However, RLHFuse goes further by eliminating the vast majority of bubbles and addressing the long-tail issue in the generation stage with intra- and inter-stage fusion. Additionally, RLHFuse integrates production-grade optimizations in the underlying framework tailored for large-scale training scenarios (§6). As a result, RLHFuse achieves 1.4×1.4$\times$1.4 ×2.4×2.4$\times$2.4 × higher throughput compared to ReaLHF. Compared to RLHFuse-Base, our system achieves a relative improvement of 1.2×1.2$\times$1.2 ×1.4×1.4$\times$1.4 × on the throughput. This improvement is entirely from stage fusion, which significantly alleviates the issues of data skewness in the generation stage and pipeline bubbles in the training stage. In the following, we conduct a detailed breakdown analysis to show the performance improvement.
@@ 7.2 Performance Analysis
For Gen.+Inf., since RLHFuse-Base does not perform inter-stage fusion and executes the two stages in serial, we use light and dark colors to represent the time spent on the generation and inference stages respectively to better show the speedup of inter-stage fusion. It can be seen that as the maximum generation length increases, the processing time of long-tailed samples is long enough for RLHFuse to fully overlap the inference stage execution, achieving 1.2×1.2$\times$1.2 ×1.6×1.6$\times$1.6 × speedup. For the training stage, similarly, we use light and dark colors to represent the Actor and Critic training time in RLHFuse-Base. Through intra-stage fusion, RLHFuse greatly mitigates pipeline bubbles and reduces the execution time of the training stage by 1.2×1.2$\times$1.2 ×1.3×1.3$\times$1.3 ×. As for the other overhead, thanks to our optimizations in task switching (§6), it only accounts for less than 3% of the total iteration time, and the migration overhead of inter-stage fusion is also negligible.
Figure 9: The fused generation and inference stage execution time of RLHFuse under different migration ratios and model settings. The maximum generation length is set to 1024.